User name M.Olson
Log entry time 08:03:06 on June 4,2007
Entry number 204220
keyword=Owl Shift Summary
Owl Shift Summary:
Kinematics Change at Beginning of Shift to Q2=68 fm^-2
Note: Could not move RHRS camera. Fortunately, we had a vernier option available that was already near the center of the screen.
0048: Kinematics change complete Q2=68 fm^-2
start run 4026 (start run dialog not updated)
0104: End run 4026 to adjust prescale factors
0106: Start run 4027 (PS1=3500, PS2=15200, PS3=1, PS4=86, PS5,6,7=1)
0138: End run 4027 (He4, Q2=68, 50/60uA, 1.25M)
0158: End run 4028 (He4, Q2=68, 50uA, 693k)
Set PS3 from 1 to 2 to keep dead time to ~2%
0159: Start run 4029 (PS3 = 2)
0258: End run 4029 (He4, Q2=68, 50uA, 1.25M)
0355: End run 4030 (He4, Q2=68, 60uA, 1.25M)
0500: End run 4031 (He4, Q2=68, 55/60uA, 1.25M)
0609: End run 4032 (He4, Q2=68, 60uA, 1.25M)
0710: End run 4033 (He4, Q2=68, 60uA, 1.25M)
0742: Current now 65uA.
Run 4034 in progress at end of shift.
Average Current (Runs 4026-4033) = 45.8 uA (from endruns)